# Reconstruction of boosted and resolved multi-Higgs-boson events with symmetry-preserving attention networks

### v1

@@ 3.2 SPA-Net
The tensor attention output is interpreted as the assignment probability ($\mathrm{AP}$). Each entry of the tensor represents the likelihood that a set of objects reconstructs the target. The indices of the tensor entry indicate which objects are to be assigned to the target. The assignment outputs are trained only on examples in which the event contains all detector objects necessary for a correct target assignment, i.e. the target is reconstructible. Nonreconstructible targets are ignored via the masked loss defined in Eq. (6) of Ref. Fenton:2020woz . As a result, the SPA-Net assignment probability only represents a conditional assignment distribution over jet indices for each target given that the target is reconstructible.
Since Ref. Fenton:2020woz introduced the ability to reconstruct partial events, it is important to estimate the probability that a given target is reconstructible in the event. This detection probability is estimated with an additional output head of SPA-Net. Specifically, a summary target vector is extracted from each of the target transformer encoders. These target vectors are fed into a feedforward binary classification network to produce a detection probability for each target. We also take into account the event-level symmetries in a similar manner to the assignment reconstruction loss from Eq. (6) of Ref. Fenton:2020woz . This detection output is trained with a cross-entropy loss over the symmetric target masks.
@@ 3.3 Target Higgs Boson Mass Reweighting
The mass distribution of the resolved and boosted Higgs boson targets in the combined $\mathrm{HHH}$ dataset is shown in Fig. 5 (left). Despite varying the Higgs boson mass between 120 and 130 GeV, both the resolved and boosted targets still possess an obvious peak around $m_{\mathrm{H}}=125\,\text{GeV}$. An overabundance of samples at $m_{\mathrm{H}}=125\,\text{GeV}$ in the training set could bias the SPA-Net algorithm to reconstruct Higgs bosons by simply choosing the jet assignment where the invariant mass is closest to $m_{\mathrm{H}}=125\,\text{GeV}$. This behavior is undesirable because a reconstruction that heavily relies on the invariant mass can result in mass sculpting of the background, that is introducing an artificial signal-like peak in the invariant mass distribution even when no signal is present.
@@
Figure 5: The mass distributions of target jet assignments show peak-like shapes, despite the Higgs targets being generated at uniformly spaced mass points (left). The weights derived to flatten the mass spectrum in the range $m\in[75,175]\,\text{GeV}$ (center). Outside the mass range $[75,175]\,\text{GeV}$, the weights are one. The resulting mass distribution after reweighting (right).
@@
To further prevent SPA-Net from reconstructing Higgs bosons only based on the invariant mass of the jet assignment, we implemented a binned reweighting as a function of the Higgs boson target mass $m\in[75,175]\,\text{GeV}$. The weights were calculated independently for each topology and were incorporated in the loss term calculation. For each topology and each mass bin of width 5 GeV, we calculated the weight as
@@
$w(m)=\begin{cases}\displaystyle\frac{p_{\mathrm{uniform}}(m_{i})}{p_{\mathrm{mass}}(m_{i})},&75\leq m_{i}\leq m<m_{i+1}\leq 175\,\text{GeV}\\ 1,&m\notin[75,175]\,\text{GeV}\end{cases},$ (3)
@@
where $p_{\mathrm{mass}}$ is the real mass distribution and $p_{\mathrm{uniform}}$ is the desired uniform distribution, and $m_{i}$ and $m_{i+1}$ are bin edges. The calculated weight for each bin for each topology is shown in Fig, 5 (center). We assigned a weight of one to the targets outside the mass range [75, 175] GeV to prevent assigning a high weight to outliers. The weights are incorporated into the calculations for both the assignment loss and the detection loss. For instance, the assignment loss is modified as follows,
$\mathcal{L}_{\mathrm{assignment}}=\min_{\sigma\in G_{t}}\sum_{i=1}^{N_{\mathrm{target}}}w_{\sigma(i)}\mathrm{CCE}(P_{i},T_{\sigma(i)}),$ (4)
where $\mathrm{CCE}$ represents the categorical cross-entropy loss, $G_{t}$ denotes the permutation group of all targets, $\sigma$ is an element of the permutation group, $N_{\mathrm{target}}$ represents the maximum number of the targets that the SPA-Net model can output, $P_{i}$ is the $i$th predicted jet assignment, $T_{\sigma(i)}$ is the $i$th target jet assignment under permutation $\sigma$, and $w_{\sigma(i)}$ is the calculated weight associated with $T_{\sigma(i)}$. Equation (4) can also be applied to the detection loss by replacing the categorical cross-entropy loss with the binary cross-entropy loss.
@@
where $[x\equiv y]$ represents a function that outputs 1 if the jet assignments $x$ and $y$ are equivalent, and 0 otherwise. Another key difference between our validation metric and the default SPA-Net validation metric lies in the selection of events used for the computation. The default SPA-Net validation accuracy is averaged only across events in which all possible target particles are reconstructible (fully reconstructible events). However, we found that we needed to include partially reconstructible events as well when incorporating boosted and resolved targets because only a small fraction of events contain three Higgs bosons that can be reconstructed in both topologies.
@@ 4 Postprocessing
We postprocess the outputs for the two topologies separately. For each topology, we calculate the most probable number of Higgs bosons in that topology based on the multinomial distribution. Here, $\mathrm{DP}_{i}$ represents the detection probability of the $i$th Higgs boson in a given topology of $\mathrm{HHH}$ events. Using a threshold of 0.5, we consider the $i$th Higgs boson to be detected if $\mathrm{DP}_{i}$ is greater than or equal to 0.5. The corresponding probabilities for detecting a specific number of Higgs bosons, $N_{\mathrm{H}}$, are then calculated as:
$p(N_{\mathrm{H}}=0)$ $=(1-\mathrm{DP}_{1})(1-\mathrm{DP}_{2})(1-\mathrm{DP}_{3})$ (6)
$p(N_{\mathrm{H}}=1)$ $=\mathrm{DP}_{1}(1-\mathrm{DP}_{2})(1-\mathrm{DP}_{3})$
$+\mathrm{DP}_{2}(1-\mathrm{DP}_{1})(1-\mathrm{DP}_{3})$
$+\mathrm{DP}_{3}(1-\mathrm{DP}_{1})(1-\mathrm{DP}_{2})$ (7)
$p(N_{\mathrm{H}}=2)$ $=\mathrm{DP}_{1}\mathrm{DP}_{2}(1-\mathrm{DP}_{3})+\mathrm{DP}_{2}\mathrm{DP}_{3}(1-\mathrm{DP}_{1})+\mathrm{DP}_{3}\mathrm{DP}_{1}(1-\mathrm{DP}_{2})$ (8)
$p(N_{\mathrm{H}}=3)$ $=\mathrm{DP}_{1}\mathrm{DP}_{2}\mathrm{DP}_{3}.$ (9)
@@
Additionally, to accurately analyze Higgs bosons originating from both resolved and boosted topologies, it is crucial to implement strategies that prevent double counting. Without such measures, the same Higgs boson could be reconstructed in both topologies and appear twice in different categories. In this study, we address this issue by prioritizing boosted Higgs boson candidates as those have a significantly better signal-to-background ratio. Specifically, resolved Higgs boson candidates are discarded if any of their associated AK5 jets overlap with boosted candidates. The overlap condition was defined using a separation criterion of $\Delta R<0.5$.
@@ 5 Results
Cosine Annealing Cycles 5 5 1 1
@@ 5.1 $\mathrm{HHH}$ Resolved Training
Figure 6 presents the Higgs reconstruction purity and efficiency as functions of the Higgs boson candidate transverse momentum $p_{\mathrm{T}}$ for the fully resolved SPA-Net training. Below $p_{\mathrm{T}}<600\,\text{GeV}$, SPA-Net demonstrates a higher reconstruction efficiency than the $\chi^{2}$ baseline, except in the first bin.
@@
As described in Section 4, we discard jet assignments where $\hat{N}_{\mathrm{H}}<3$, which indicates that not all Higgs bosons could be reconstructed. This procedure aims to balance reconstruction purity and efficiency and results in enhanced reconstruction purity with respect to the baseline, as shown in Fig. 6. However, this balancing strategy can occasionally discard valid jet assignments, reducing the reconstruction efficiency. For analyses prioritizing reconstruction efficiency, the $\mathrm{DP}$ threshold (effectively set to 0.5 in this study) could be lowered to retain more jet assignments, albeit at the cost of reduced reconstruction purity.
The jet assignment performance is affected by the angular separation of the two jets from the Higgs boson decay, which depends on the Higgs boson momentum. For low $p_{\mathrm{T}}$, the jets are widely separated and distributed across the detector, making the jet assignment challenging for three Higgs bosons. For high $p_{\mathrm{T}}$, the jets are more collimated and the jet correlation is more evident. However, above $p_{\mathrm{T}}>400\,\text{GeV}$, the efficiency drops due to the increased probability of merging two jets into a large-radius jet or losing one of the $\mathrm{b}$ jets out of the detector acceptance. This limitation can be overcome by incorporating the large-radius AK8 jets in the SPA-Net training.
@@ 5.2 $\mathrm{HHH}$ Boosted+Resolved Training
Figure 7 shows the Higgs reconstruction purity and efficiency versus the Higgs boson candidate transverse momentum $p_{\mathrm{T}}$ for the SPA-Net model trained on both resolved and boosted targets. Compared with Fig. 6, the boosted+resolved SPA-Net exhibits significantly better performance above $p_{\mathrm{T}}>400\,\text{GeV}$ as a result of including boosted targtes in the training.
@@
As introduced in Section 3.1, the baseline utilizes the $\chi^{2}$ method to assign AK5 jets to resolved Higgs boson targets and applies the BDT to classify AK8 jets into boosted Higgs boson candidates or background. Similar to SPA-Net, which predicts and reconstructs candidates in both topologies for each event, the $\chi^{2}$-based resolved baseline and the BDT-based boosted baseline also operate on each event. In this study, we assume the $\chi^{2}$ baseline always predicted three resolved Higgs boson candidates in the resolved topology. For the boosted topology, we select the AK8 jets of which BDT scores passing the loose working point of 0.911348, corresponding to a background misidentification rate of 2%. Finally, the Higgs boson candidates predicted by both baselines undergo the same overlap-removal procedure in Section 4.
@@ 5.3 $\mathrm{HH}$ Resolved and Boosted+Resolved Trainings
As $\mathrm{HH}\to 4b$ events are more likely to be detected at the LHC, we trained one SPA-Net using $\mathrm{HH}$ resolved targets and another using both $\mathrm{HH}$ boosted and resolved boosted targets. We then evaluate their Higgs boson reconstruction purity and efficiency. The corresponding Higgs boson reconstruction results are shown in Figs. 9 and 10, respectively.
@@
Similar to the $\mathrm{HHH}$ results, the $\mathrm{HH}$ results demonstrate that SPA-Net achieves higher Higgs boson reconstruction purity and efficiency than the baseline, modified to target two reconstructed Higgs bosons with the first four jets ranked in Section 3.1, across all $p_{\mathrm{T}}$ bins. Furthermore, the comparison between the resolved SPA-Net and the boosted+resolved SPA-Net shows that incorporating boosted Higgs boson targets can enhance SPA-Net’s Higgs boson reconstruction purity and efficiency for $p_{\mathrm{T}}>400\,\text{GeV}$.

### v3

@@ 3.2 SPA-Net
The tensor attention output is interpreted as the assignment probability ($\mathrm{AP}$). Each entry of the tensor represents the likelihood that a set of objects reconstructs the target. The indices of the tensor entry indicate which objects are to be assigned to the target. To train SPA-Net to output the assignment tensor, we adopt the combined symmetric loss from the original SPA-Net Shmakov:2021qdz and disable the class-balance term, as we found it degraded performance on the majority of events. Explicitly, the assignment loss is
$\mathcal{L}_{\mathrm{assignment}}=\min_{\sigma\in G_{t}}\sum_{i=1}^{N_{\mathrm{target}}}\mathcal{M}_{\sigma(i)}\mathrm{CCE}(P_{i},T_{\sigma(i)}),$ (2)
where $\mathrm{CCE}$ represents the categorical cross-entropy loss, $G_{t}$ denotes the permutation group of all targets, restricted to permutations within the same reconstruction topology, $\sigma$ is an element of the permutation group, $N_{\mathrm{target}}$ represents the maximum number of the targets that the SPA-Net model can output, $P_{i}$ is the $i$th predicted jet assignment, $T_{\sigma(i)}$ is the $i$th target jet assignment under permutation $\sigma$, and $\mathcal{M}_{\sigma(i)}$ is the target mask associated with $T_{\sigma(i)}$. The assignment outputs are trained only on examples in which the event contains all detector objects necessary for a correct target assignment, i.e. the target is reconstructible. Nonreconstructible targets are ignored via the mask $\mathcal{M}_{\sigma(i)}$ in Eq. (2). As a result, the SPA-Net assignment probability only represents a conditional assignment distribution over jet indices for each target given that the target is reconstructible.
Since Ref. Fenton:2020woz introduced the ability to reconstruct partial events, it is important to estimate the probability that a given target is reconstructible in the event. This detection probability (DP) is estimated with an additional output head of SPA-Net. We also take into account the event-level symmetries in a similar manner to the assignment loss. Specifically, a summary target vector is extracted from each of the target transformer encoders. These target vectors are fed into a feedforward binary classification network to produce a detection probability for each target. This target detection output is trained with the detection loss,
$\mathcal{L}_{\mathrm{detection}}=\min_{\sigma\in G_{t}}\sum_{i=1}^{N_{\mathrm{target}}}\mathrm{BCE}(\mathrm{DP}_{i},\mathcal{M}_{\sigma(i)}),$ (3)
where BCE denotes the binary cross-entropy loss, and $\mathrm{DP}_{i}$ is the predicted detection probability of the $i$th target.
@@ 3.3 Target Higgs Boson Mass Reweighting
The mass distribution of the resolved and boosted Higgs boson targets in the combined $\mathrm{HHH}$ dataset is shown in Fig. 7 (left). Despite varying the Higgs boson mass between 120 and 130 GeV, both the resolved and boosted targets still possess an obvious peak around $m_{\mathrm{H}}=125\,\text{Ge\hskip-0.80002ptV}$. An overabundance of samples at $m_{\mathrm{H}}=125\,\text{Ge\hskip-0.80002ptV}$ in the training set could bias the SPA-Net algorithm to reconstruct Higgs bosons by simply choosing the jet assignment where the invariant mass is closest to $m_{\mathrm{H}}=125\,\text{Ge\hskip-0.80002ptV}$. This behavior is undesirable because a reconstruction that heavily relies on the invariant mass can result in mass sculpting of the background, that is introducing an artificial signal-like peak in the invariant mass distribution even when no signal is present.
@@
Figure 7: The mass distributions of target jet assignments show peak-like shapes, despite the Higgs targets being generated at uniformly spaced mass points (left). The weights derived to flatten the mass spectrum in the range $m\in[75,175]\,\text{Ge\hskip-0.74ptV}$ (center). Outside the mass range $75,175]\,\text{Ge\hskip-0.74ptV}$, the weights are one. The resulting mass distribution after reweighting (right).
@@
To further prevent SPA-Net from reconstructing Higgs bosons only based on the invariant mass of the jet assignment, we implemented a binned reweighting as a function of the Higgs boson target mass $m\in[75,175]\,\text{Ge\hskip-0.80002ptV}$. The weights were calculated independently for each topology and were incorporated in the loss term calculation. For each topology and each mass bin of width 5 GeV, we calculated the weight as
@@
$w(m)=\begin{cases}\displaystyle\frac{p_{\mathrm{uniform}}(m_{i})}{p_{\mathrm{mass}}(m_{i})},&75\leq m_{i}\leq m<m_{i+1}\leq 175\,\text{Ge\hskip-0.80002ptV}\\ 1,&m\notin[75,175]\,\text{Ge\hskip-0.80002ptV}\end{cases},$ (5)
@@
where $p_{\mathrm{mass}}$ is the real mass distribution and $p_{\mathrm{uniform}}$ is the desired uniform distribution, and $m_{i}$ and $m_{i+1}$ are bin edges. The calculated weight for each bin for each topology is shown in Fig. 7 (center). We assigned a weight of one to the targets outside the mass range [75, 175] GeV to prevent assigning a high weight to outliers. The weights are incorporated into the calculations for both the assignment loss and the detection loss. For instance, the assignment loss for each event is modified to be
$\mathcal{L}_{\mathrm{assignment}}=\min_{\sigma\in G_{t}}\sum_{i=1}^{N_{\mathrm{target}}}w_{\sigma(i)}\mathcal{M}_{\sigma(i)}\mathrm{CCE}(P_{i},T_{\sigma(i)}),$ (6)
where, in contrast to Eq. (2), the assignment loss term for each target is scaled by the corresponding target-wise weight $w_{\sigma(i)}$. Similarly, the detection loss is modified by multiplying the loss term for each target in the event by $w_{\sigma(i)}$ .
@@
where $x\equiv y]$ represents a function that outputs 1 if the jet assignments $x$ and $y$ are equivalent, and 0 otherwise. Another key difference between our validation metric and the default SPA-Net validation metric lies in the selection of events used for the computation. The default SPA-Net validation accuracy is averaged only across events in which all possible target particles are reconstructible (fully reconstructible events). However, we found that we needed to include partially reconstructible events as well when incorporating boosted and resolved targets because only a small fraction of events contain three Higgs bosons that can be reconstructed in both topologies.
@@ 4 Postprocessing
We postprocess the outputs for the two topologies separately. For each topology, we calculate the most probable number of Higgs bosons in that topology based on the multinomial distribution. The probabilities for detecting a specific number of Higgs bosons, $N_{\mathrm{H}}$, are calculated as:
$p(N_{\mathrm{H}})$ $=\begin{cases}(1-\mathrm{DP}_{1})(1-\mathrm{DP}_{2})(1-\mathrm{DP}_{3})&N_{\mathrm{H}}=0\\ \mathrm{DP}_{1}(1-\mathrm{DP}_{2})(1-\mathrm{DP}_{3})+\mathrm{DP}_{2}(1-\mathrm{DP}_{1})(1-\mathrm{DP}_{3})&N_{\mathrm{H}}=1\\ {}+\mathrm{DP}_{3}(1-\mathrm{DP}_{1})(1-\mathrm{DP}_{2})\\ \mathrm{DP}_{1}\mathrm{DP}_{2}(1-\mathrm{DP}_{3})+\mathrm{DP}_{2}\mathrm{DP}_{3}(1-\mathrm{DP}_{1})+\mathrm{DP}_{3}\mathrm{DP}_{1}(1-\mathrm{DP}_{2})&N_{\mathrm{H}}=2\\ \mathrm{DP}_{1}\mathrm{DP}_{2}\mathrm{DP}_{3}&N_{\mathrm{H}}=3\end{cases}.$ (8)
@@
Additionally, to accurately analyze Higgs bosons originating from both resolved and boosted topologies, it is crucial to implement strategies that prevent double counting. Without such measures, the same Higgs boson could be reconstructed in both topologies and appear twice in different categories. In this study, we address this issue by prioritizing boosted Higgs boson candidates, as these typically have a significantly better signal-to-background ratio Butterworth:2008iy ; CMS:2017bcq . Specifically, resolved Higgs boson candidates are discarded if any of their associated AK5 jets overlap with an AK8 jet corresponding to a boosted Higgs boson candidate. The overlap condition is defined using a separation criterion of $\Delta R<0.5$. Similarly, for the purpose of uniquely identifying a (boosted or resolved) Higgs boson target for post-training evaluation metrics, resolved Higgs boson targets are discarded if the same Higgs boson also matches an AK8 jet according to the criteria in Section 2.
@@ 5 Results
Cosine Annealing Cycles 5 1 1 1
@@ 5.1 $\mathrm{HHH}$ Resolved Training
Figure 8 presents the Higgs reconstruction purity and efficiency as functions of the Higgs boson candidate transverse momentum $p_{\mathrm{T}}$ for the fully resolved SPA-Net training. Below $p_{\mathrm{T}}<600\,\text{Ge\hskip-0.80002ptV}$, SPA-Net demonstrates a higher reconstruction efficiency than the $\chi^{2}$ baseline, except in the first bin.
@@
As described in Section 4, we discard jet assignments where $\hat{N}_{\mathrm{H}}<3$, which indicates that not all Higgs bosons could be reconstructed. This procedure aims to balance reconstruction purity and efficiency and results in enhanced reconstruction purity with respect to the baseline, as shown in Fig. 8. However, this balancing strategy can occasionally discard valid jet assignments, reducing the reconstruction efficiency.
The jet assignment performance is affected by the angular separation of the two jets from the Higgs boson decay, which depends on the Higgs boson momentum. For low $p_{\mathrm{T}}$, the jets are widely separated and distributed across the detector, making the jet assignment challenging for three Higgs bosons. For high $p_{\mathrm{T}}$, the jets are more collimated and the jet correlation is more evident. However, above $p_{\mathrm{T}}>400\,\text{Ge\hskip-0.80002ptV}$, the efficiency drops due to the increased probability of merging two jets into a large-radius jet or losing one of the $\mathrm{b}$ jets out of the detector acceptance. This limitation can be overcome by incorporating the large-radius AK8 jets in the SPA-Net training.
@@ 5.2 $\mathrm{HHH}$ Boosted+Resolved Training
Figure 9 shows the Higgs reconstruction purity and efficiency versus the Higgs boson candidate transverse momentum $p_{\mathrm{T}}$ for the SPA-Net model trained on both resolved and boosted targets. Compared with Fig. 8, the boosted+resolved SPA-Net exhibits significantly better performance above $p_{\mathrm{T}}>400\,\text{Ge\hskip-0.80002ptV}$ as a result of including boosted targets in the training.
@@
As introduced in Section 3.1, the baseline utilizes the $\chi^{2}$ method to assign AK5 jets to resolved Higgs boson targets and applies the BDT to classify AK8 jets into boosted Higgs boson candidates or background. Similar to SPA-Net, which predicts and reconstructs candidates in both topologies for each event, the $\chi^{2}$-based resolved baseline and the BDT-based boosted baseline also operate on each event. In this study, we assume the $\chi^{2}$ baseline always predicted three resolved Higgs boson candidates in the resolved topology. For the boosted topology, we select the AK8 jets of which BDT scores passing the loose working point of 0.911348, corresponding to a signal efficiency of 38% and a background misidentification rate of 2%, as shown in Fig. 5. Finally, the Higgs boson candidates predicted by both baselines undergo the same overlap-removal procedure in Section 4.
@@
Figure 10: Confusion matrix for the full events in the testing fraction of the $\mathrm{HHH}$ $m_{\mathrm{H}}=125\,\text{Ge\hskip-0.74ptV}$ dataset. The true label indicates the number of boosted and resolved-only targets in each event, while the predicted label corresponds to the number of boosted and resolved Higgs boson candidates identified after postprocessing. Most true categories maximized along the diagonal, indicating strong event topology classification performance.
Additionally, we present the confusion matrix for events with three reconstructible Higgs bosons target (referred to as full events) in Fig. 10. An event is considered full if all true Higgs bosons can be successfully matched to either 2 AK5 jets or 1 AK8 jet as described in Section 2. We focus on full events as they are expected to dominate the signal sensitivity in experimental searches. The classification categories are as follows: three resolved target Higgs bosons (0bh3rh), one boosted and two resolved targets (1bh2rh), two boosted and one resolved target (2bh1rh), and three boosted targets (3bh0rh). For each event, the true category is determined by counting the number of boosted targets and the number of resolved targets that do not correspond to any boosted target. The predicted category is obtained by counting the number of boosted candidates and resolved candidates after applying the postprocessing procedure described in Section 4. The matrix is normalized over the true category (rows).
Most true categories are predicted correctly, indicating the strong event topology classification performance of SPA-Net and our postprocessing method. An exception is the 3bh0rh category, which is underrepresented in the dataset (0.1%), leading to reduced prediction accuracy.
@@ 5.3 $\mathrm{HH}$ Resolved and Boosted+Resolved Trainings
As $\mathrm{HH}\to 4b$ events are more likely to be detected at the LHC, we trained one SPA-Net using $\mathrm{HH}$ resolved targets and another using both $\mathrm{HH}$ boosted and resolved boosted targets. We then evaluate their Higgs boson reconstruction purity and efficiency. The corresponding Higgs boson reconstruction results are shown in Figs. 12 and 13, respectively.
@@
Similar to the $\mathrm{HHH}$ results, the $\mathrm{HH}$ results demonstrate that SPA-Net achieves higher Higgs boson reconstruction purity and efficiency than the baseline, modified to target two reconstructed Higgs bosons with the first four jets ranked in Section 3.1, across all $p_{\mathrm{T}}$ bins. Furthermore, the comparison between the resolved SPA-Net and the boosted+resolved SPA-Net shows that incorporating boosted Higgs boson targets can enhance SPA-Net’s Higgs boson reconstruction purity and efficiency for $p_{\mathrm{T}}>400\,\text{Ge\hskip-0.80002ptV}$.
Figure 14: Confusion matrix for full events in the testing fraction of the $\mathrm{HH}$ $m_{\mathrm{H}}=125\,\text{Ge\hskip-0.74ptV}$ dataset. Compared to Fig. 10, all true categories are maximized along the diagonal, indicating strong event topology classification performance.
Similar to the confusion matrix in Fig. 10, Fig. 14 presents the confusion matrix for full events in the $\mathrm{HH}$ dataset, illustrating that SPA-Net, combined with our postprocessing method, accurately classifies each target category based on the number of resolved and boosted Higgs boson candidates.